OpenStudy (anonymous):

An angle has a measure of (5x + 7) and another angle has a measure of (2x + 54). Together they form a linear pair. Find the measure of the smaller angle.

OpenStudy (raden):

5x+7+2x+54 = 180 solve for x, first

OpenStudy (anonymous):

So after combining like terms I got 68 so I would then divide 68 by itself and 180, correct?

OpenStudy (raden):

after u combining the similar term, that should be 5x+7+2x+54 = 180 7x + 61 = 180 subtract by 61 to both sides, giving us 7x + 61 - 61 = 180 - 61 7x = 119 divide by 7 to both sides, 7x/7 = 119/7 x = 17

OpenStudy (anonymous):

Oh okay I see where I had my mistake, thank you very much. :)

OpenStudy (raden):

now, subtitute x = 17 to the measure 5x+7 and (2x+54) respectively to comparing whose as the smaller angle then ...

OpenStudy (anonymous):

And the smaller angle would be 88, which was from the (2x+54) angle.